f(x) =10x^3-3x^2-5x-1


1
Expert's answer
2021-07-12T14:03:53-0400

Maclaurin series of function "f(x)" is defined as


"f(x)=f(0)+\dfrac{f'(0)}{1!}x+\dfrac{f''(0)}{2!}x^2+\dfrac{f'''(0)}{3!}x^3+\dfrac{f''''(0)}{4!}x^4+..."


"f(x) =10x^3-3x^2-5x-1"


"f(0)=-1"


"f'(x)=30x^2-6x-5, f'(0)=-5"


"f''(x)=60x-6, f''(0)=-6"


"f'''(x)=60, f'''(0)=60"


"f^{(IV)}(x)=f^{(V)}(x)=f^{(VI)}(x)=...=0"


Then


"f(x)=-1+\dfrac{-5}{1}x+\dfrac{-6}{2}x^2+\dfrac{60}{6}x^3"

"c_0=-1, c_1=-5, c_2=-3, c_4=10, c_5=c_6=...=0"

"f(x)=-1-5x-3x^2+10x^3"

The radius of convergence "R=INF."